Docstoc

FGC Software Design Review - LHC commissioning_2_

Document Sample
FGC Software Design Review - LHC commissioning_2_ Powered By Docstoc
					        Wednesday 13th February
    n 06:58 Beams dumped by OP, ~2 pb-1 in 8h18m
    n 09:01 Stable beams
             ¨ Initial lumi 145e30 cm-2s-1, ATLAS had a DAQ problem and did
               not take data for the first part of the fill
    n 10:30 Train moving in the tunnel
    n 14:08 Beams dumped, 1.6 pb-1 in 5h07m
             ¨ SIS interlock on orbit tolerances when starting Length Scale
               calib at pt 1. SIS COD tolerances for 1.38TeV were too tight.
    n 18:57 Stable beams
             ¨ Test COD for V/d Meer scans
             ¨ V/d Meer scans ATLAS “done”… almost
    n 21:03 Beams dumped by BLM spurious trigger
             ¨ Also IP1, as on Tuesday, but different card

                                               Many telephone calls later….
14/02/2013                             LHC 8:30 meeting
        pp 1.38 TeV run



             2 pb-1




                      500 b
                              1374 b
                                                          RP !

                                                   V/dM


              MON         Tue                Wed


14/02/2013                LHC 8:30 meeting
        The incredible night (extending the extension)
    n Laurette and Markus going for it
             ¨ 23:00 JJ in for BCT calibration
             ¨ 23:45 Stable Beams, Checking the CODs
             ¨ 00:08 Start CMS V/d Meer scan
             ¨ 01:24 CMS length scale
             ¨ 02:14 ATLAS ready for length scale…
             ¨ 03:36 Repeating VdM scan in separation plane !
             ¨ 04:17 Moving Roman Pots in at 13 sigma. Loss maps: 4 x
             ¨ 04:38 Beams dumped
                n Belen checking loss maps: OK
                n JJ again for reconfiguration of BCTs.
             ¨ 06:37 Stable Beams – turn around < 2 h !
                n RP in at 13 sigma, taking data
                n SMS from Stefano, call from Mario, …
             ¨ 07:24 Last dump from Stable Beams for a long time…
                n Operator dump!
14/02/2013                               LHC 8:30 meeting
             End of Run 1 (2009 - 2013)




14/02/2013           LHC 8:30 meeting
        Train                                          (Bruno Feral)
     n 10:30 Start moving the train with beam
     n Up to now the train is fully operational and controllable from surface
       with stable vision feedback.
     n On the RP side we had a maximum of 1mSv when exiting UL55 with
       an average of 0.3µSv along the machine.
     n Arrive point 4
     n 15:03 train @ garage




14/02/2013                         LHC 8:30 meeting
        It is not over yet…
     n 48 h Quench tests + 3h pc limits IPQs
        ¨EPC, 3 hours during working hours




14/02/2013                   LHC 8:30 meeting

				
DOCUMENT INFO
Shared By:
Categories:
Tags:
Stats:
views:5
posted:11/11/2013
language:
pages:6